# Neural Spectro-polarimetric Fields

### v1

@@ 7. Discussion
NeSpoF interprets light as rays, thus does not encompass phase and amplitude of light [Goodman and Sutton, 1996]. Also, NeSpoF demands hours of training time and tuning parameters that may benefit from recently-proposed neural representations. On the hardware front, our imaging system requires a lengthy capture time of 10 seconds for a single view. Developing a single-shot hyperspectral-polarimetric camera leaves as a future direction. Lastly, there is potential for further exploring the applications of NeSpoF in uncovering geometry, appearance, and material properties.
@@ 8. Conclusion
We propose NeSpoF, the first neural representation specifically designed for hyperspectral Stokes vector fields. Demonstrated across synthetic and real-world scenes captured via a calibrated hyperspectral-polarimetric imager, NeSpoF reveals hidden spectro-polarimetric-positional information of the scene outside the confines of human vision. We anticipate that NeSpoF spurs further interest in high-dimensional imaging, modeling, and analysis.

### v2

@@ 7. Discussion
While NeSpoF aims to model additional wave properties of light, spectrum and polarization, NeSpoF still interprets light as rays, thus does not encompass phase and amplitude of light [Goodman and Sutton, 1996]. Also, NeSpoF demands hours of training time and tuning parameters that may benefit from recently-proposed neural representations. On the hardware front, our imaging system requires a lengthy capture time of 10 seconds for a single view. Also, the captured images contain considerable measurement noise due to the low transmission of the employed optical filters. We leave developing a single-shot light-efficient hyperspectral-polarimetric camera as a future direction. Lastly, there is potential for further exploring the applications of NeSpoF in uncovering geometry, appearance, and material properties, similar to the progress on the applications of RGB radiance fields.
@@ 8. Conclusion
We propose NeSpoF, the first neural representation specifically designed for physically-valid hyperspectral Stokes-vector fields. Demonstrated across synthetic scenes and real-world scenes captured via a calibrated hyperspectral-polarimetric imager, NeSpoF reveals hidden spectro-polarimetric-positional information of the scene outside the confines of human vision. We anticipate that NeSpoF spurs further interest in high-dimensional imaging, modeling, and analysis.
Acknowledgements.
This work was partly supported by Korea NRF (RS-2023-00211658, 2022R1A6A1A03052954, 2023R1A2C200494611), Samsung Research Funding Center (SRFCIT1801-52), Samsung Electronics, Korea MOTIE (NTIS1415187366, 20025752), and Korea IITP MSIT (No.2019-0-01906, Artificial Intelligence Graduate School Program-POSTECH).